# Textual Similarity as a Key Metric in Machine Translation Quality Estimation

### v1

@@ Abstract
Machine Translation (MT) Quality Estimation (QE) assesses translation reliability without reference texts. This study introduces “textual similarity” as a new metric for QE, using sentence transformers and cosine similarity to measure semantic closeness. Analyzing data from the MLQE-PE dataset, we found that textual similarity exhibits stronger correlations with human scores than traditional metrics (hter, model evaluation etc.). Employing GAMMs as a statistical tool, we demonstrated that textual similarity consistently outperforms other metrics across multiple language pairs in predicting human scores. We also found that “hter” actually failed to predict human scores in QE. Our findings highlight the effectiveness of textual similarity as a robust QE metric, recommending its integration with other metrics into QE frameworks and MT system training for improved accuracy and usability.
@@ 1 Introduction
However, in MT QE, several features play pivotal roles in both quality estimation and enhancing the overall translation process. The “model eveluation” (ML_eval) quantifies the MT model’s confidence in its translations. This metric is typically derived from the log-likelihood of the translation given the source sentence. Higher scores reflect greater confidence, indicating that the translation aligns well with the model’s learned patterns from the training data. These scores are crucial for quality estimation TM systems, helping to predict the reliability of translations from TM models and highlight areas that may require further human intervention or correction [2], [4].
@@
While “ML_eval” provides valuable insights into QE, the metric has certain limitations. One significant issue is the tendency to overestimate confidence in poor translations, often due to overfitting or a lack of diverse training data. This overconfidence can lead to inaccuracies in QE. Moreover, the metric is sensitive to the training data distribution, and meaning biases or gaps in the data can skew the scores. Third, this metric may also fail to capture contextual meaning and subtle semantic differences that human evaluators can detect, leading to discrepancies between model scores and perceived translation quality.
“hter” has also some weaknesses. First, the variability in human editing styles and preferences can lead to inconsistent hter scores, making it challenging to standardize quality assessments. Additionally, “hter” focuses on the number of edits rather than the nature of the changes, failing to differentiate between minor stylistic tweaks and substantial corrections. This can skew the perception of translation quality. Furthermore, the metric relies on the availability of high-quality reference translations, which may not always be accessible, limiting its applicability in certain contexts. These limitations suggest that while “hter” is useful, it should be complemented with other evaluation methods for a more comprehensive assessment of translation quality.
@@ 2.2 Textual similarity
Textual similarity (or sentence similarity) measures how similar or different two pieces of text are semantically. This can involve comparing sentences, paragraphs, or entire documents to see how closely they match in meaning or content. It can be calculated using various techniques, such as comparing word overlaps, semantic similarity, or using machine learning models to assess how alike the texts are. HHowever, after transformers revolutionized deep learning, textual similarity can now be effectively measured using existing language models. [12], [3], [13]
@@
The “sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2” model creates dense vector representations of sentences, capturing the semantic essence of the text. These vectors are placed in a shared embedding space where similar meanings are mapped to nearby points. The model’s fine-tuning for tasks like paraphrase identification ensures that sentences with similar meanings are close in this space. The use of sentence transformers is particularly valuable in multilingual contexts, as they support multiple languages and capture semantic nuances effectively. The task in the current study is to understand the source text in one language and the translated texts in the other language. The capability of multilingual sentence transformer models is essential for applications to explore the semantic similarity among languages. The current study employs this multilingual sentence transformer model to compute semantic similarity for one source sentence and its translation. Our fundamental approach involves using the sentence transformer to generate text embeddings and then calculating their similarity using the cosine method.
@@ 3.1 Result 1: Correlations
We plotted the Pearson correlations among various factors, as shown in Fig. LABEL:fig. The results indicate that “ML_eval” is correlated with the human score (mean) at 0.15, with “hter” at 0.06, and with the human score (z-mean) at 0.3. “Textual similarity” shows a correlation of 0.47 with the human score (mean), -0.06 with the human score (z-mean), -0.45 with “hter”, and -0.21 with “ML_eval”. Overall, “textual similarity” exhibits stronger correlations with other factors compared to “ML_eval”.
Figure 1: correlations among various factors
3.2 Result 2: Overall performance
We established three types of GAMM fittings, where “human score (mean)” is the independent variable, and other factors include “ML_eval”, “textual similarity”, “sd of human scores”. “human evaluator number”, “language pairs (langs)” are incorporated as random variables. We chose “human score (mean)” over “human score (z-mean)” as the independent variable primarily because “human score (mean)” closely follows a normal distribution. The GAMM setups are detailed below, and the results are presented in Table 1.
@@ 3.2 Result 2: Overall performance
Here, the base model demonstrates the best performance when all factors are included. However, each of the remaining models includes only a subset of these factors. The Akaike Information Criterion (AIC) is used to represent the performance of this GAMM fitting. The base model has the lowest AIC. When the AIC of model $m1$ is subtracted from the AIC of the base model, the resulting $\Delta$AIC indicates the contribution of “textual similarity”, as “m1” does not include “textual similarity” compared to the base model. A smaller $\Delta$AIC also indicates better performance and greater contribution for a given factor. The results are shown in Table 1
Table 1: $\Delta$AIC for different GAMM fittings with random smooths. A smaller $\Delta$AIC indicates better performance (n=45886)
@@
The model with “textual similarity” outperforms the one with “ML_eval”. However, the model with “hter” shows the best performance overall. This is an overall trend. Despite this, it is important to examine the performance across individual language pairs. The next section explores the performance of these factors in each language pair.
3.3 Result 3: Across language pairs
Using the similar GAMM setups in the previous section, we explore how these factors perform across seven language pairs. For each language pair, we established the same GAMM fittings (i.e., the base model, the model without textual similarity, the model without “ML_eval”, and the model without “hter” ). $\Delta$AIC is the subtraction of a specific model from the base model. Note that the random variable “ language pairs” is not included because we implmented GAMM fittings for an indiviual language pair among the seven pairs. After carrying out four GAMM fittings for each of seven language pairs, we plotted the partial effects for each factor of interest, as shown in Figure 2. When the p-value is greater than 0.05, the plot is not significant. Within the same language pair, a smaller $\Delta$AIC indicates better performance. Clearly, all cases for “ML_eval” are significant, and there is only one insignificant case for “textual similarity”. In contrast, “hter” has only two significant cases, indicating five insignificant cases.
@@ 3.3 Result 3: Across language pairs
Figure 2: The partial effects on human score from different factors. The x-axis represents the specific metric being analyzed, while the y-axis indicates human score. Each curve within a plot illustrates the relationship between a predictor variable (plotted on the x-axis) and the response variable. Steeper slopes on these curves indicate a stronger influence of the predictor variable on human score. Conversely, gentler slopes imply a weaker influence, indicating that changes in the predictor variable have a less pronounced effect on human score. Such plots could give deep insights on the relationship between one given metric and human score. The information on p-value and $\Delta$AIC is shown at the top of each plot.
@@ 4 Discussion
Considering the correlation, we found that “textual similarity” exhibits a strong correlation with human scores, “ML_eval”, and “hter”. In contrast, “ML_eval” shows a weaker correlation with the human score (mean). When examining the GAMM fittings overall and across various language pairs, the performance of “textual similarity” surpasses that of “ML_eval”. On an individual language pair basis, “textual similarity” also outperforms both “ML_eval” and “hter”. In short, “textual similarity” as a metric could be closely correlated with human score, and is highly capable of predicting human score.
The following provides a detailed analysis of each metric. “ML_eval” demonstrates a significant impact across all language pairs, indicating that this metric is both useful and effective for QE. However, “hter” does not show a significant impact on human scores in the most cases, suggesting that this metric may not be suitable for evaluating MT quality in certain language pairs. On the other hand, “textual similarity” predicts human scores in the majority of language pairs, highlighting its potential as an effective metric for QE.
@@
Therefore, it is worthwhile to explore the cognitive recognition of what constitutes a good translation for different language users. Understanding this can significantly enhance both machine translation and quality estimation processes by aligning them more closely with human judgments of translation quality.
To date, “textual similarity” has not been proposed as a metric in QE. Our study applied a variety of statistical methods to analyze data from established QE datasets, demonstrating that this metric is both reliable and effective compared to commonly used metrics in QE. Our findings provide strong evidence that “textual similarity” is a robust metric, making it a valuable addition to the existing suite of QE metrics. Given its proven effectiveness, “textual similarity” should be included as a key metric in QE and incorporated into the training of machine translation (MT) systems. This integration can enhance the accuracy and reliability of MT quality assessments, ultimately improving the performance and usability of MT systems.

### v2

@@ Abstract
Machine Translation (MT) Quality Estimation (QE) assesses translation reliability without reference texts. This study introduces “textual similarity” as a new metric for QE, using sentence transformers and cosine similarity to measure semantic closeness. Analyzing data from the MLQE-PE dataset, we found that textual similarity exhibits stronger correlations with human scores than traditional metrics (hter, model evaluation, sentence probability etc.). Employing GAMMs as a statistical tool, we demonstrated that textual similarity consistently outperforms other metrics across multiple language pairs in predicting human scores. We also found that “hter” actually failed to predict human scores in QE. Our findings highlight the effectiveness of textual similarity as a robust QE metric, recommending its integration with other metrics into QE frameworks and MT system training for improved accuracy and usability.
@@ 1 Introduction
However, in MT QE, several features play pivotal roles in both quality estimation and enhancing the overall translation process. The “model eveluation” (ML_eval) quantifies the MT model’s confidence in its translations. This metric is typically derived from the log-likelihood of the translation given the source sentence. Higher scores reflect greater confidence, indicating that the translation aligns well with the model’s learned patterns from the training data. These scores are crucial for quality estimation TM systems, helping to predict the reliability of translations from TM models and highlight areas that may require further human intervention or correction [2], [6]. Moreover, in many cases, sentences with higher probabilities from n-gram models do tend to be more natural-sounding. This is because these models are trained on large corpora of text, capturing common patterns in language use. Sentence probabilities for translated texts are also used to evaluate translation quality without reference translations.
@@
While “ML_eval” provides valuable insights into QE, the metric has certain limitations. One significant issue is the tendency to overestimate confidence in poor translations, often due to overfitting or a lack of diverse training data. This overconfidence can lead to inaccuracies in QE. Moreover, the metric is sensitive to the training data distribution, and meaning biases or gaps in the data can skew the scores. Third, this metric may also fail to capture contextual meaning and subtle semantic differences that human evaluators can detect, leading to discrepancies between model scores and perceived translation quality. Additionally, although higher probability from an n-gram model often correlates with more natural-sounding sentences, this metric may not be a ideal indicator. The reason for this is that the relationship between statistical probability and perceived naturalness is complex and influenced by many factors beyond n-gram probabilities.
The metric, “hter”, has also some weaknesses. First, the variability in human editing styles and preferences can lead to inconsistent hter scores, making it challenging to standardize quality assessments. Additionally, “hter” focuses on the number of edits rather than the nature of the changes, failing to differentiate between minor stylistic tweaks and substantial corrections. This can skew the perception of translation quality. Furthermore, the metric relies on the availability of high-quality reference translations, which may not always be accessible, limiting its applicability in certain contexts. These limitations suggest that while “hter” is useful, it should be complemented with other evaluation methods for a more comprehensive assessment of translation quality.
@@ 2.1 Dataset
This study utilized two distinct datasets. The first dataset, MLQE-PE [6], and the second dataset, used in PreQuEL [5]. Although the datasets differ in size, they can both be effectively employed for cross-validation purposes.
@@
PreQuEL [5] includes a variety of datasets. However, we selected only the datasets for evaluating machine translation quality without reference translations. In this case, there are two language pairs: English-Chinese and English-German. The dataset includes the following variables: n-gram sentence probability, language model score, HTER, and human score (z_mean). “Sentence probability” refers to the likelihood of a particular sequence of words (sentence) occurring. This can be computed using n-grams. For example, if bi-grams are used, it is termed “bi-gram sentence probability”, and if tri-grams are used, it is “tri-gram sentence probability”. PreQuEL provides five types of sentence probabilities using 1-5 grams. The “language model score” is a metric related to the language probability score (“lan” is indicated in the PreQuEL). This score is part of the quality estimation process that evaluates how likely a translation is accurate based on linguistic features and models, without referencing the actual translated text. PreQuEL also provides “hter”, which is computed using the same method as in MLQE-PE. Additionally, PreQuEL provides only the “human score (z_mean)”.
@@ 2.2 Textual similarity
Textual similarity (or sentence similarity) measures how similar or different two pieces of text are semantically. This can involve comparing sentences, paragraphs, or entire documents to see how closely they match in meaning or content. It can be calculated using various techniques, such as comparing word overlaps, semantic similarity, or using machine learning models to assess how alike the texts are. However, after transformers revolutionized deep learning, textual similarity can now be effectively measured using existing language models. [14], [4], [15]
@@
The “sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2” model creates dense vector representations of sentences, capturing the semantic essence of the text. These vectors are placed in a shared embedding space where similar meanings are mapped to nearby points. The model’s fine-tuning for tasks like paraphrase identification ensures that sentences with similar meanings are close in this space. The use of sentence transformers is particularly valuable in multilingual contexts, as they support multiple languages and capture semantic nuances effectively. The task in the current study is to understand the source text in one language and the translated texts in the other language. The capability of multilingual sentence transformer models is essential for applications to explore the semantic similarity among languages. The current study employs this multilingual sentence transformer model to compute semantic similarity for one source sentence and its translation. Our fundamental approach involves using the sentence transformer to generate text embeddings and then calculating their similarity using the cosine method. We applied the method to process the two datasets to compute “textual similarity” between a source text and its translated text.
@@ 3.1 Result 1: Correlations
We plotted the Pearson correlations among various factors for MLQE-PE, as shown in Fig. 1. The results indicate that “ML_eval” is correlated with the human score (mean) at 0.15, with “hter” at 0.06, and with the human score (z-mean) at 0.3. “Textual similarity” shows a correlation of 0.47 with the human score (mean), -0.06 with the human score (z-mean), -0.45 with “hter”, and -0.21 with “ML_eval”. Overall, “textual similarity” exhibits stronger correlations with other factors compared to “ML_eval”.
Figure 1: correlations among various factors in MLQE-PE
Similarly, we plotted the correlation among various factors for PreQuEL, as shown in Fig 4 (see the Appendix). The three variables: “n-gram sentence probability”, “hter”, and “textual similarity”, show similar correlation values with the “human score (z_mean)”, each approximately 0.15.
3.2 Result 2: Cross language pairs
We established three types of GAMM fittings for MLQE-PE, where “human score (mean)” is the independent variable, and other factors include “ML_eval”, “textual similarity”, “sd of human scores”. “human evaluator number”, “language pairs (langs)” are incorporated as random variables. We chose “human score (mean)” over “human score (z-mean)” as the independent variable primarily because “human score (mean)” closely follows a normal distribution. The GAMM setups are detailed below, and the results are presented in Table 1.
@@ 3.2 Result 2: Cross language pairs
The base model demonstrates the best performance when all factors are included. However, each of the remaining models includes only a subset of these factors. The Akaike Information Criterion (AIC) is used to represent the performance of this GAMM fitting. The base model has the lowest AIC. When the AIC of model $m1$ is subtracted from the AIC of the base model, the resulting $\Delta$AIC indicates the contribution of “textual similarity”, as “m1” does not include “textual similarity” compared to the base model. A smaller $\Delta$AIC also indicates better performance and greater contribution for a given factor. The results are shown in Table 1.
Table 1: $\Delta$AIC for different GAMM fittings for MLQE-PE. A smaller $\Delta$AIC indicates better performance (n=45886)
@@
The contribution from “textual similarity” outperforms that from “ML_eval”. However, the contribution of “hter” shows the best performance overall. This trend holds for all language pairs. Despite this, it is important to examine the performance for individual language pairs. The next section explores the performance of these factors in each language pair.
Next, we used similar GAMM methods to explore how different metrics affect human scores for PreQuEL. In the PreQuEL dataset, “human score (z_mean)” is the dependent variable  [^2] In this dataset, “human score (mean)” is not available., and other factors include “n-gram sentence probability”, “language model score”, “hter”, and “textual similarity”. In this dataset, “language model score” (“lm_score”) has four values, so we treat it as a random factor, and “language pairs” (“langs”) are incorporated as a random variable. There are five types of “n-gram sentence probability”; however, we chose the optimal one, “trigram sentence probability”,"for the GAMM fittings. The GAMM setups are detailed below, and the results are presented in Table 2.
  • base model = bam(human score $\sim$ s(trigram sent prob)+s(textual
    similarity)+s(hter) +s(lm_score, bs=“re”)+s(langs, bs=“re”))
  • t1=bam(human score $\sim$ s(trigram sent prob)+s(hter)+s(lm_score, bs=“re”)
    +s(langs, bs=“re”))
  • t2=bam(human score $\sim$ s(textual similarity)+s(hter)+s(lm_score, bs=“re”)
    +s(langs, bs=“re”))
  • t3=bam(human score $\sim$ s(trigram sent prob)+s(textual similarity)
    +s(lm_score, bs=“re”)+s(langs, bs=“re”))
Table 2: $\Delta$AIC for different GAMM fittings for PreQuEL. A smaller $\Delta$AIC indicates better performance (n=14706)
factor (contribution) $\Delta$AIC
(t1- base model) contribution of textual similarity 237.11
(t2- base model) contribution of trigram sentence probability 232.16
(t3-base model) contribution of hter 424.29
The contribution of “textual similarity” outperforms that of “hter”. However, the contribution of “trigram sentence probability” shows a similar performance to “textual similarity”, with “trigram sentence probability” demonstrating the best performance overall. The next section examines the performance for individual language pairs.
3.3 Result 3: Individual language pair
Using a similar GAMM setup as the base model, we explore how these factors perform across seven language pairs for MLQE-PE. For each language pair, we established the same GAMM and plotted the partial effects for each factor of interest, as shown in Figure 2. When the p-value is greater than 0.05, the plot is not significant. Within the same language pair, a smaller $\Delta$AIC indicates better performance. Clearly, all cases for “ML_eval” are significant, and there is only one insignificant case for “textual similarity”. In contrast, “hter” has only two significant cases, indicating five insignificant cases.
@@ 3.3 Result 3: Individual language pair
Figure 2: The partial effects on human score from different factors for MLQE-PE. The x-axis represents the specific metric being analyzed, while the y-axis indicates human score. Each curve within a plot illustrates the relationship between a predictor variable (plotted on the x-axis) and the response variable. Steeper slopes on these curves indicate a stronger influence of the predictor variable on human score. Conversely, gentler slopes imply a weaker influence, indicating that changes in the predictor variable have a less pronounced effect on human score. Such plots could give deep insights on the relationship between one given metric and human score.
Adopting the similar GAMM fittings for PreQuEL, we explored the performance for each variable of our interest in each language pair, and plotted the partial effects for each factor of interest, as shown in Figure 3. Figure 3 shows that “ textual similariy” has the best performance in each individual language pair.
Figure 3: The partial effects on human score from different factors for PreQuEL. The x-axis represents the specific metric being analyzed, while the y-axis indicates the human score. The interpretation of the curve is the same as in Fig 2. The layout here differs slightly from that in Fig 2. $\Delta$AIC values are compared among three plots for the same language pairs. A lower $\Delta$AIC value indicates better performance.
@@ 4 Discussion
Considering the correlation in MLQE-PE, we found that “textual similarity” exhibits a strong correlation with human scores, “ML_eval”, and “hter”. In contrast, “ML_eval” shows a weaker correlation with the human score (mean). When examining the GAMM fittings across various language pairs and in each case of language pair, the performance of “textual similarity” surpasses that of “ML_eval”. On an individual language pair basis, “textual similarity” also outperforms both “ML_eval” and “hter”. In PreQuEL, due to the smaller sample size, the differences in correlation are not obvious. In short, “textual similarity” as a metric could be closely correlated with human score, and is highly capable of predicting human score.
The following provides a detailed analysis of each metric in MLQE-PE. “ML_eval” demonstrates a significant impact across all language pairs, indicating that this metric is both useful and effective for QE. However, “hter” does not show a significant impact on human scores in the most cases, suggesting that this metric may not be suitable for evaluating MT quality in certain language pairs. On the other hand, “textual similarity” predicts human scores in the majority of language pairs, highlighting its potential as an effective metric for QE.
@@
Next, we analyzed the performance of “textual similarity” in PreQuEL. “textual similarity” outperforms “hter” consistently across language and within invidiual languages. However, “textual similarity” has the similar performance with “n-gram sentence probability”. It also reveals that “n-gram sentence probability” may be a useful metric in evaluating MT QE.
@@
Therefore, exploring the cognitive recognition of what constitutes a good translation for different language users is worthwhile. Understanding this can significantly enhance both machine translation and quality estimation processes by aligning them more closely with human judgments of translation quality. For instance, "n-gram sentence probability" can provide an initial impression of the naturalness or readability of translated texts. If this metric is low, human evaluators might rate the quality of the translations poorly, regardless of how closely they match the original meaning. Simple yet effective metrics should be considered in evaluating machine translation quality, incorporating factors that align with human translation assessment standards and processes. This approach can lead to substantial progress in machine translation quality estimation (MT QE).
To date, “textual similarity” has not been proposed as a metric in QE  [^3] Although some studies have proposed using semantic similarity to evaluate MT quality, they have only applied this method in cases where reference translations are provided [3, 17]. Specifically, these studies compared the semantic similarity between a candidate translation and a reference translation. In contrast, our study employed cross-lingual the semantic similarity between a text in the source language and its translated text in the target language.. Our study applied a variety of statistical methods to analyze data from established QE datasets, demonstrating that this metric is both reliable and effective compared to commonly used metrics in QE. Our findings provide strong evidence that “textual similarity” is a robust metric, making it a valuable addition to the existing suite of QE metrics. Given its proven effectiveness, “textual similarity” should be included as a key metric in QE and incorporated into the training of machine translation (MT) systems. This integration can enhance the accuracy and reliability of MT quality assessments, ultimately improving the performance and usability of MT systems.
Appendix
Figure 4: correlations among various factors in PreQuEL